# Codes Correcting Two Bursts of Exactly $b$ Deletions

### v3

@@ 1. $(\textup{{P}}1)$
Suppose $\mathbf{y}$ is obtained from a codeword $\mathbf{x}\in\mathcal{C}$ by two bursts of erasures. Then $A(\mathbf{y})_{i}$ is obtained from $A(\mathbf{x})_{i}$ by two erasures for each $i=1,2$. Suppose the error positions in $A(\mathbf{y})_{1}$ are $i_{1},i_{2}$, and the error positions in $A(\mathbf{y})_{2}$ are $i_{3},i_{4}$, where $i_{1}<i_{2}$ and $i_{3}\leq i_{4}$. Clearly, we have $i_{3}\in\{i_{1},i_{1}-1\}$ and $i_{4}\in\{i_{2},i_{2}-1\}$. This implies that $i_{4}-i_{3}\in\{i_{2}-i_{1},i_{2}-i_{1}-1,i_{2}-i_{1}+1\}$. By definition, the two bursts in $\mathbf{x}$ do not overlap. It follows that $i_{2}\geq i_{1}+2$ if $i_{4}-i_{3}=i_{2}-i_{1}-1$.
To decode $\mathbf{x}$ from $\mathbf{y}$. We first calculate the two erased symbols in each row. Notice that $N>2$ and hence $Q$ is an odd prime. Suppose $\sigma_{1},\sigma_{2}\in[0,N-1]$ satisfy
$\sigma_{1}+\sigma_{2}\equiv t_{1}\pmod{Q},$ (1)
$\sigma_{1}^{2}+\sigma_{2}^{2}\equiv t_{2}\pmod{Q}$
where $t_{1}$ and $t_{2}$ are given. Then we can uniquely recover the (multi)set $\{\sigma_{1},\sigma_{2}\}$ by solving system (1). Now we can determine the erased symbols in $A(\mathbf{y})_{1}$ and $A(\mathbf{y})_{2}$ by (C1). Suppose that the two erased symbols in $A(\mathbf{y})_{1}$ are $\alpha_{1},\alpha_{2}$, and the two erased symbols in $A(\mathbf{y})_{2}$ are $\alpha_{3},\alpha_{3}$. It remains to determine positions of $\alpha_{1}$ and $\alpha_{2}$ in $A(\mathbf{x})_{1}$, and positions of $\alpha_{3}$ and $\alpha_{4}$ in $A(\mathbf{x})_{2}$. We can assume $\alpha_{1}\neq\alpha_{2}$ or $\alpha_{3}\neq\alpha_{4}$, since otherwise, we can uniquely recover $\mathbf{x}$.
There are at most four choices for the positions of $\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}$ and for each choice, we get a candidate for $\mathbf{x}$. We need to show that among these candidates, there is exactly one equals $\mathbf{x}$. Firstly, by the relationship of $\mathbf{x}$ and $\mathbf{y}$, it is clear that there is at least candidate equals $\mathbf{x}$. Next, we will show there is at most one such candidate. To that end, let $(\beta_{1},\beta_{2})$ be a permutation of $(\alpha_{1},\alpha_{2})$, and $(\beta_{3},\beta_{4})$ be a permutation of $(\alpha_{3},\alpha_{4})$, where $(\beta_{1},\beta_{2})\neq(\alpha_{1},\alpha_{2})$ or $(\beta_{3},\beta_{4})\neq(\alpha_{3},\alpha_{4})$. Let $\mathbf{u}$ be obtained from $\mathbf{y}$ by assigning $\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}$ at coordinates $i_{1},i_{2},i_{3},i_{4}$, respectively. Let $\mathbf{v}$ be obtained from $\mathbf{y}$ by assigning $\beta_{1},\beta_{2},\beta_{3},\beta_{4}$ at coordinates $i_{1},i_{2},i_{3},i_{4}$, respectively. By assumption, we have $\mathbf{u}\neq\mathbf{v}$. We need to show that at least one of $\mathbf{u}$ and $\mathbf{v}$ is not a codeword in $\mathcal{C}$. Let $\Delta=W(\mathbf{u})-W(\mathbf{v})$. Then we have
$\Delta=i_{1}(\alpha_{1}-\beta_{1})+i_{2}(\alpha_{2}-\beta_{2})+2N\left(i_{3}(\alpha_{3}-\beta_{3})+i_{4}(\alpha_{4}-\beta_{4})\right).$
To complete the proof, it is sufficient to show that at least one of $\mathbf{u}$ and $\mathbf{v}$ does not satisfies (C2). Suppose on the contrary that $W(\mathbf{u})\equiv c\equiv W(\mathbf{v})\pmod{nN^{2}}$. Then we have $\Delta\equiv 0\pmod{nN^{2}}$.
If $(\beta_{1},\beta_{2})=(\alpha_{1},\alpha_{2})$, it holds that $(\beta_{3},\beta_{4})\neq(\alpha_{3},\alpha_{4})$. Therefore, we conclude that $\alpha_{3}\neq\alpha_{4}$ and $(\beta_{3},\beta_{4})=(\alpha_{4},\alpha_{3})$. Then it follow that $\left|\Delta\right|=2N(i_{4}-i_{3})\left|\alpha_{3}-\alpha_{4}\right|$ and hence $0<\left|\Delta\right|\leq\left\lfloor\frac{n}{2}\right\rfloor 2N(N-1)<nN^{2}$, which is a contradiction. Similarly, if $(\beta_{3},\beta_{4})=(\alpha_{3},\alpha_{4})$, we have $0<\left|\Delta\right|\leq\left\lfloor\frac{n}{2}\right\rfloor(N-1)<nN^{2}$, which is a contradiction.
Now suppose $\alpha_{1}\neq\alpha_{2}$, $\alpha_{3}\neq\alpha_{4}$, $(\beta_{1},\beta_{2})=(\alpha_{2},\alpha_{1})$ and $(\beta_{3},\beta_{4})=(\alpha_{4},\alpha_{3})$. Then we have $\Delta=(i_{2}-i_{1})(\alpha_{2}-\alpha_{1})+2N(i_{4}-i_{3})(\alpha_{4}-\alpha_{3})$. Without loss of generality, suppose $\alpha_{1}<\alpha_{2}$. If $\alpha_{3}<\alpha_{4}$, then $0<\Delta\leq\left\lfloor\frac{n}{2}\right\rfloor(N-1)+\left\lfloor\frac{n}{2}\right\rfloor 2N(N-1)\leq n(N^{2}-1)$, which is a contradiction. Now suppose $\alpha_{3}>\alpha_{4}$. Recall that $i_{4}-i_{3}\in\{i_{2}-i_{1}-1,i_{2}-i_{1},i_{2}-i_{1}+1\}$. When $i_{4}-i_{3}=i_{2}-i_{1}$ or $i_{2}-i_{1}+1$, it is easy to verify that $0<-\Delta\leq\left\lfloor\frac{n}{2}\right\rfloor(2N(N-1)-1)+2N(N-1)<nN^{2}$, which is a contradiction. When $i_{4}-i_{3}=i_{2}-i_{1}-1$, we have $i_{2}\geq i_{1}+2$. Then it follows that
$0<2N(\alpha_{3}-\alpha_{4})-2(\alpha_{2}-\alpha_{1})\leq-\Delta\leq\left\lfloor\frac{n}{2}\right\rfloor\left(2N(N-1)-1\right)-2N<nN^{2},$
which is a contradiction. ∎
@@ III Bounds
We construct a graph $G$ where the vertex set $V(G)$ is $\Sigma_{q}^{n}$ and two distinct vertices $\mathbf{x}$, $\mathbf{y}$ in $\Sigma_{q}^{n}$ is connected by an edge (denoted by $\mathbf{x}\sim\mathbf{y}$) if and only if $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})\neq\emptyset$. An independent set of $G$ is a subset of $\Sigma_{q}^{n}$ such that any two distinct vertices are not connected by an edge. Let $\alpha(G)$ be the maximum size of an independent set of $G$. It is clear that a subset $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is a $2$-$b$-burst-deletion correcting code if and only if $\mathcal{C}$ is an independent set in $G$. Therefore, we have $M_{q,n,b}=\alpha(G)$. For a vertex $\mathbf{x}$, let $d(\mathbf{x})$ be the number of $\mathbf{y}$ such that $\mathbf{x}\sim\mathbf{y}$. Then it follows from [37, page 100, Theorem 1] that
@@
In [38, Section IV-B], the authors proved an upper bound of $1$-$b$-burst-deletion correcting codes. Next, we adapt their idea to derive an upper bound on $M_{q,n,b}$. Recall that $\ln(\cdot)$ is the natural logarithm function.
@@
where (a) follows from the well-known result (see the proof of [40, Theorem 3.1]):
@@
Since $t=(1-1/q-\epsilon)m$, we have $t/m=1-1/q-\epsilon\leq 1-1/q$. Then it follows from [41, Proposition 3.3.3] that
@@
If $\frac{\log n}{n}<\frac{\log q}{12}f(q)^{2}$, then $\epsilon<f(q)$. So by Lemma .1, we have.
@@
Now it follows from Equations 5 and 6 that
@@
where (a) follows from selecting $\epsilon=\sqrt{\frac{12\log n}{n\log q}}$ and (b) follows from the fact that $n]{n}<1.121$ when $n\geq 30$. We conclude the proof by combining Equations 4 and 7. ∎
@@
IV Binary $2$-$b$-Burst-Deletion Correcting Codes
In this section, we will construct a class of binary $2$-$b$-burst-deletion correcting codes with redundancy at most $7\log n+O(\log\log n)$. First, we introduce some $1$-deletion correcting codes and $2$-deletion correcting codes, which will be useful later.
Lemma IV.1
[21, Lemma 1] For any integer $n\geq 3$, there exists a function $\mathsf{VT}:\Sigma_{2}^{n}\rightarrow\Sigma_{2}^{\log(n+1)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{2}^{n}$, given $\mathsf{VT}\left(\mathbf{x}\right)$ and any $\mathbf{y}\in\mathcal{B}_{1}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$.
@@ IV Binary $2$-$b$-Burst-Deletion Correcting Codes
Lemma IV.4
[21, Lemma 9] Suppose that $\mathbf{x}\in\Sigma^{n}$ is $d$-regular and $\mathbf{x}^{\prime}\in\Sigma^{n-2}$ is obtained from $\mathbf{x}$ by deleting two symbols $x_{i_{1}}$ and $x_{i_{2}}$. Then we can

### v4

@@ 1. $(\textup{{P}}1)$
Suppose $\mathbf{y}$ is obtained from a codeword $\mathbf{x}\in\mathcal{C}$ by two bursts of erasures. Then $A(\mathbf{y})_{i}$ is obtained from $A(\mathbf{x})_{i}$ by two erasures for each $i=1,2$. Denote the error positions in $A(\mathbf{y})_{1}$ are $i_{1},i_{2}$, and the error positions in $A(\mathbf{y})_{2}$ are $i_{3},i_{4}$, where $i_{1}<i_{2}$ and $i_{3}\leq i_{4}$. Note that $i_{1},i_{2},i_{3}$ and $i_{4}$ are known to us. Clearly, we have $i_{3}\in\{i_{1},i_{1}-1\}$ and $i_{4}\in\{i_{2},i_{2}-1\}$. This implies that $i_{4}-i_{3}\in\{i_{2}-i_{1},i_{2}-i_{1}-1,i_{2}-i_{1}+1\}$.
Next, we describe how to decode $\mathbf{x}$ from $\mathbf{y}$. Clearly, it is sufficient to recover the values of $A(\mathbf{x})_{1,i_{1}}$, $A(\mathbf{x})_{1,i_{2}}$, $A(\mathbf{x})_{2,i_{3}}$ and $A(\mathbf{x})_{2,i_{4}}$. Let
$\delta_{1}=\left(a_{1}-\sum_{j=1,j\neq i_{1},i_{2}}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{1,j}\right)\pmod{2N},$ (1)
$\delta_{2}=\left(a_{2}-\sum_{j=1,j\neq i_{3},i_{4}}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{2,j}\right)\pmod{2N}.$
Since $0\leq A(\mathbf{x})_{1,i_{1}}+A(\mathbf{x})_{1,i_{2}},A(\mathbf{x})_{2,i_{3}}+A(\mathbf{x})_{2,i_{4}}<2N$ and $0\leq\delta_{1},\delta_{2}<2N$, it follows from Condition (C1) and Equation 1 that $\delta_{1}=A(\mathbf{x})_{1,i_{1}}+A(\mathbf{x})_{1,i_{2}}$ and $\delta_{2}=A(\mathbf{x})_{2,i_{3}}+A(\mathbf{x})_{2,i_{4}}$.
For simpler notations, denote $\alpha_{1}=A(\mathbf{x})_{1,i_{1}}$ and $\alpha_{2}=A(\mathbf{x})_{2,i_{3}}$. Then we have $A(\mathbf{x})_{1,i_{2}}=\delta_{1}-\alpha_{1}$ and $A(\mathbf{x})_{2,i_{4}}=\delta_{2}-\alpha_{2}$. Therefore, it remains to obtain the values of $\alpha_{1}$ and $\alpha_{2}$. To that end, let
$\Delta=\left(b-\sum_{j=1,j\neq i_{1},i_{2}}^{\left\lceil n/2\right\rceil}j\cdot A(\mathbf{y})_{1,j}-(2N-1)\sum_{j=1,j\neq i_{3},i_{4}}^{\left\lceil n/2\right\rceil}j\cdot A(\mathbf{y})_{2,j}\right)\pmod{nN^{2}}.$ (2)
Since each term in the right-hand side of (2) is known, we can obtain the value of $\Delta$. Furthermore, by (2), Condition (C2), and the relationship between $A(\mathbf{x})$ and $A(\mathbf{y})$, it holds that
$\Delta$ $=\left(W(\mathbf{x})-\sum_{j=1,j\neq i_{1},i_{2}}^{\left\lceil n/2\right\rceil}j\cdot A(\mathbf{x})_{1,j}-(2N-1)\sum_{j=1,j\neq i_{3},i_{4}}^{\left\lceil n/2\right\rceil}j\cdot A(\mathbf{x})_{2,j}\right)\pmod{nN^{2}}$ (3)
$=\left(i_{1}\alpha_{1}+i_{2}(\delta_{1}-\alpha_{1})+(2N-1)(i_{3}\alpha_{2}+i_{4}(\delta_{2}-\alpha_{2}))\right)\pmod{nN^{2}}.$
Let $\Delta^{\prime}=\left(i_{2}\delta_{1}+(2N-1)i_{4}\delta_{2}-\Delta\right)\pmod{nN^{2}}$, which can be calculated since values of $\delta_{1}$, $\delta_{2}$ and $\Delta$ are known to us. Then it follows from Equation 3 that
$(i_{2}-i_{1})\alpha_{1}+(2N-1)(i_{4}-i_{3})\alpha_{2}\equiv\Delta^{\prime}\pmod{nN^{2}}.$ (4)
Since $i_{1}<i_{2}\leq\left\lceil n/2\right\rceil$ and $i_{4}-i_{3}\leq\left\lceil n/2\right\rceil$, we have $21\leq i_{2}-i_{1},i_{4}-i_{3}\leq n/$ 2. Combining this with the fact $10\leq\alpha_{1},\alpha_{2}\leq N-$ 1, we conclude that $0\leq(i_{2}-i_{1})\alpha_{1}+(2N-1)(i_{4}-i_{3})\alpha_{2}<nN^{2}$. Now Equation 4 implies
$(i_{2}-i_{1})\alpha_{1}+(2N-1)(i_{4}-i_{3})\alpha_{2}=\Delta^{\prime}.$ (5)
Recall that $i_{4}-i_{3}\in\{i_{2}-i_{1}-1,i_{2}-i_{1},i_{2}-i_{1}+1\}$. By definition, the two bursts of erasures in $\mathbf{x}$ do not overlap. It follows that $i_{2}\geq i_{1}+2$ if $i_{4}-i_{3}=i_{2}-i_{1}-1$. Therefore, we always have $i_{4}-i_{3}\geq 1$. This implies $(2N-1)(i_{4}-i_{3})>1$. Then it follows from Equation 5 that
$(i_{2}-i_{1})\alpha_{1}\equiv\Delta^{\prime}\pmod{(2N-1)(i_{4}-i_{3})}.$ (6)
Note that $10\leq\alpha_{1}\leq N-$ 1. When $i_{4}-i_{3}=i_{2}-i_{1}$ or $i_{2}-i_{1}+1$, it it easy to see that $0\leq(i_{2}-i_{1})\alpha_{1}<(2N-1)(i_{4}-i_{3})$. When $i_{4}-i_{3}=i_{2}-i_{1}-1$, since $i_{2}-i_{1}\geq 2$, we have $(2N-1)(i_{4}-i_{3})-(i_{2}-i_{1})\alpha_{1}=(i_{2}-i_{1})(2N-1-\alpha_{1})-(2N-1)\geq 2N-1-2\alpha_{1}>0$ and hence $(i_{2}-i_{1})\alpha_{1}<(2N-1)(i_{4}-i_{3})$. Now it follows from Equation 6 that $(i_{2}-i_{1})\alpha_{1}=\Delta^{\prime\prime}$, where $\Delta^{\prime\prime}=\Delta^{\prime}\pmod{(2N-1)(i_{4}-i_{3})}$. From this, we get $\alpha_{1}=\frac{\Delta^{\prime\prime}}{(i_{2}-i_{1})}$. Then by Equation 5, we have $\alpha_{2}=\frac{\Delta^{\prime}-(i_{2}-i_{1})\alpha_{1}}{(2N-1)(i_{4}-i_{3})}$. Now the proof is completed. ∎
@@ III Bounds
We construct a graph $G$ where the vertex set $V(G)$ is $\Sigma_{q}^{n}$ and two distinct vertices $\mathbf{x}$, $\mathbf{y}$ in $\Sigma_{q}^{n}$ is connected by an edge (denoted by $\mathbf{x}\sim\mathbf{y}$) if and only if $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})\neq\emptyset$. An independent set of $G$ is a subset of $\Sigma_{q}^{n}$ such that any two distinct vertices are not connected by an edge. Let $\alpha(G)$ be the maximum size of an independent set of $G$. A subset $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is a $2$-$b$-burst-deletion correcting code if and only if $\mathcal{C}$ is an independent set in $G$. Therefore, we have $M_{q,n,b}=\alpha(G)$. For a vertex $\mathbf{x}$, let $d(\mathbf{x})$ be the number of $\mathbf{y}$ such that $\mathbf{x}\sim\mathbf{y}$. Then it follows from [38, page 100, Theorem 1] that
@@
In [39, Section IV-B], the authors proved an upper bound of $1$-$b$-burst-deletion correcting codes. Next, we adapt their idea to derive an upper bound on $M_{q,n,b}$. Recall that $\ln(\cdot)$ is the natural logarithm function.
@@
where (a) follows from the well-known result (see the proof of [41, Theorem 3.1]):
@@
Since $t=(1-1/q-\epsilon)m$, we have $t/m=1-1/q-\epsilon\leq 1-1/q$. Then it follows from [42, Proposition 3.3.3] that
@@
If $\frac{\log n}{n}<\frac{\log q}{12}f(q)^{2}$, then $\epsilon<f(q)$. Now by Lemma .1, we have
@@
Now it follows from Equations 10 and 11 that
@@
where (a) follows from selecting $\epsilon=\sqrt{\frac{12\log n}{n\log q}}$ and (b) follows from the fact that $n]{n}<1.121$ when $n\geq 30$. We conclude the proof by combining Equations 9 and 12. ∎
@@
IV Codes For Correcting Two $b$-Burst-Deletions
In this section, we construct $q$-ary $2$-$b$-burst-deletion correcting codes with redundancy at most $5\log n+O(\log\log n)$, for any $q\geq 2$ and $b>1$. Our idea is to first locate positions of deletions in short intervals (which is accomplished in Section IV-A, Lemma IV.5) and then correct errors in these intervals (which is accomplished in Sections IV-B and IV-C).
IV-A Approximately determine positions of deletions
@@
Lemma IV.3
[21, Lemma 9] Suppose that $\mathbf{x}\in\Sigma_{2}^{n}$ is $d$-regular and $\mathbf{x}^{\prime}\in\Sigma_{2}^{n-2}$ is obtained from $\mathbf{x}$ by deleting two symbols $x_{i_{1}}$ and $x_{i_{2}}$. Then we can